# citations.bib

### v1

@@ I Introduction and Background
Organizations heavily rely upon third-party software to help them achieve their business objectives. For example, countless organizations use Microsoft Windows in their laptop, desktop, and server environments [RN71]. Microsoft provides these organizations with compiled binaries, but does not provide the actual source code. Due to the loss of high-level structures during compilation from languages such as C/C++, identifying vulnerabilities in compiled binaries is significantly more resource-intensive than in source code. Thus, these organizations must trust the third party (e.g., Microsoft) to follow secure coding practices when developing their applications. Unfortunately, like all software, coding mistakes are made, introducing bugs that threat actors exploit to cause mass destruction to organizations around the world. For example, the WannaCry [RN75] cryptoworm exploited an SMB vulnerability in Windows systems, encrypted the data on the infected systems, and held the data on these systems for ransom. Recovering from these types of attacks has cost organizations billions of dollars [RN76], resulted in economic inflation [ransomeinfl], and even cost innocent people their lives [hospitalfatalities].
To address the challenge of identifying vulnerabilities within compiled code, some researchers have investigated using machine learning to detect vulnerabilities within compiled code. For example, [RN9] and [schaad2022deeplearningbased] compile code samples from the NIST Juliet [RN89] dataset, lift the compiled code to LLVM using the RetDec [RN22] tool, and preprocess the lifted code. This preprocessed code was used to train a word2vec model for creating token-level embeddings, and these embeddings were provided to recurrent neural networks to learn to differentiate between vulnerable and non-vulnerable code. However, although word2vec is capable of learning high-level code semantics [mikolov2013efficient], it lacks the context awareness present in transformer-based Natural Language Processing (NLP) models [VaswaniSPUJGKP17]. [mccully1] expanded on this research by examining the application of the BERT and RoBERTa models (both bidirectional) to generate embeddings for the lifted code. These embeddings were then used to train neural networks to identify stack-based buffer overflows (CWE-121) within the lifted code. The research introduced in this paper builds on the foundational work of [RN9, schaad2022deeplearningbased, mccully1] by comparing and contrasting the impacts of using a unidirectional transformer (GPT-2), bidirectional transformers (BERT, RoBERTa), and non-transformer-based embedding models (Skip-Gram, Continuous Bag of Words) to train an LSTM vulnerability classifier.
@@ II Methodology and Implementation
Figure 1: Building GPT-2 Model Using Lifted Code
@@
Figure 2: Training Neural Network to Detect Buffer Overflows
@@ II-A Embedding Dataset Creation
The dataset used to train the GPT-2 model was generated by selecting code samples from the NIST SARD Juliet source code dataset without CWE-121 and compiling them on a Kali Linux system [^1] Originally, the samples were also compiled on a Windows 10 system. However, after inspecting the code lifted to LLVM, it was found that RetDec was stripping calls to the C standard library. Thus, these samples were removed from the dataset.. The Juliet dataset contains code samples for several vulnerability classes and the corresponding code samples that have implemented a fix to remediate the vulnerability. The code samples were turned into object files on Linux by running the ‘make’ Juliet1.3 command.
@@ II-B Neural Network Dataset Creation
A subset of the Juliet dataset was used to create the dataset used to train a neural network to differentiate between LLVM functions that contain buffer overflow vulnerabilities and those that do not. As noted previously, the Juliet dataset contains code samples for several vulnerabilities including CWE-121 (Stack-Based Buffer Overflow). For the purposes of this research, these code samples were compiled into object files, and the functions were extracted from these files. The Juliet dataset uses a standard naming scheme that indicates functions that contain buffer overflows and those in which the vulnerability was mitigated. The preprocessing steps in II-C were applied to the data set, but after the preprocessing was completed, it was discovered that some of the functions contained minimal functionality, such as calling a secondary function that contained or did not contain the vulnerability. After preprocessing, most of these functions contained identical instructions and were removed from the dataset when duplicates were eliminated. Finally, functions greater than 2048 tokens in length were removed from the dataset due to resource constraints on neural network training systems. The breakdown of the size of the CWE-121 dataset is given in Table I. Given the relatively small dataset, rather than balancing it with an equal number of vulnerable and non-vulnerable functions, metrics such as precision, recall, and F1-score were utilized to evaluate the overall model performance.
@@ II-C LLVM Preprocessing
The preprocessing phase removes content unique to each function without affecting how the function works. For example, function names are unique to each function, but do not affect how the LLVM code executes. Thus, a function name of UniqueFunctionName can be standardized to Func_one without losing functionality. The purpose of preprocessing is to remove these aspects of the LLVM functions to standardize how the functions look. The rationale for this choice is the belief that excessive superfluous information within the LLVM functions could hinder the models’ ability to learn the most relevant code-level patterns. This study followed the preprocessing steps described in detail in [mccully1].
@@ II-D GPT-2 Model Creation
For this study, the GPT2LMHeadModel HuggingFace library was used to create a GPT-2 model that was trained using the same LLVM functions used to train the word2vec, BERT, and RoBERTa models in  [mccully1]. The system used to train the model contained dual TI 4090 GPUs. The tokenizer used to tokenize the LLVM functions was the ByteLevelBPETokenizer tokenizer, and the maximum size of the LLVM functions was set to 2048 tokens. The number of hidden layers was set to 12, the dimensionality of the embeddings was set to 100, and the number of attention heads was set to 10. Finally, the embedding dataset detailed in Section II-A was split into two subsets; one subdataset for training that contained 90% of the overall dataset, and a second for validation. The model ran for 20 epochs and steadily improved accuracy every 1,000 steps, as demonstrated in Table II.
@@ IV Discussion
IV-A Unidirectional vs. Bidirectional Embedding Models
@@
IV-B Unidirectional vs. word2vec Models
The top-performing neural network trained with GPT-2 embeddings surpassed the performance of the best model that used word2vec CBOW and Skip-Gram embeddings. This is unsurprising since word2vec models contain a single hidden layer used to determine the relationship among the surrounding tokens [mikolov2013efficient]. These models lack context-awareness, meaning that, once trained, a token is always assigned the same value despite its sequential location and the tokens surrounding it. In contrast, GPT-2 is a transformer-based model capable of learning complex semantic relationships [radford2019language]. GPT-2 considers the position and surrounding tokens when assigning values to specific tokens.
@@ V-A Dataset Sizes
V-B LLVM Function Size
Due to memory constraints on the systems used to train LSTM models using GPT-2 embeddings, researchers set the maximum size of LLVM functions that could be provided to train and evaluate LSTM neural networks to 2048 tokens. This meant that approximately 100 samples from the LSTM Training Dataset had to be removed. More work would be needed to determine the best way to represent functions with more than 2048 tokens without placing too great a strain on server resources.
V-C Data Snooping
@@
*arp2022and [arp2022and] states that data snooping is one of the biggest pitfalls in research on the application of machine learning to problems in the cybersecurity field. Data snooping occurs when training and validation data are not strictly segmented. In the context of the current study, data snooping was not a factor since the samples containing CWE-121 were not used to train the GPT-2 model. However, due to the prominence of the issue within existing research [arp2022and], the effects of data snooping on model accuracy when embedding models are trained with samples in the neural network validation set will be studied in greater detail in future work.
V-D Embedding Layer Tuning Enhancements
@@ VI Related Work
Although several papers have employed LLMs to identify vulnerabilities in source code, considerably fewer have focused on using them to detect specific categories of vulnerabilities (e.g. buffer overflows) in compiled code. In [RN67], researchers train a RoBERTa model from scratch using assembly code instructions. The embeddings from this model are used to train a Message Passing Neural Network (MPNN) to differentiate between vulnerable and non-vulnerable code. [gallagherllvm] also built a RoBERTa model to learn code-level semantics to train a neural network to detect vulnerabilities. However, rather than training the RoBERTa model with assembly code, the researchers trained it with LLVM code. [han2022binary] trained a BERT model using pcode; these pcode embeddings were used to train a model that could detect vulnerabilities within the lifted code.

### v2

@@ I Introduction and Background
Organizations heavily rely upon third-party software to help them achieve their business objectives. For example, countless organizations use Microsoft Windows in their laptop, desktop, and server environments [RN71]. Microsoft provides these organizations with compiled binaries but does not provide the actual source code. Due to the loss of high-level structures during compilation from languages such as C/C++, identifying vulnerabilities in compiled binaries is significantly more resource-intensive than in source code. Thus, these organizations must trust the third party (e.g., Microsoft) to follow secure coding practices when developing their applications. Unfortunately, like all software, coding mistakes are made, introducing bugs that threat actors exploit to cause mass destruction to organizations around the world. For example, the WannaCry [RN75] cryptoworm exploited an SMB vulnerability in Windows systems, encrypted the data on the infected systems, and held the data on these systems for ransom. Recovering from these types of attacks has cost organizations billions of dollars [RN76], resulted in economic inflation [ransomeinfl], and even cost innocent people their lives [hospitalfatalities].
To address the challenge of identifying vulnerabilities within compiled code, some researchers have investigated using machine learning to detect vulnerabilities within compiled code. For example, [RN9] and [schaad2022deeplearningbased] compiled code samples from the NIST Juliet [RN89] dataset, lifted the compiled code to LLVM using the RetDec [RN22] tool, and preprocessed the lifted code. This preprocessed code was used to train a word2vec model for creating token-level embeddings, and these embeddings were provided to recurrent neural networks to learn to differentiate between vulnerable and non-vulnerable code. However, although word2vec is capable of learning high-level code semantics [mikolov2013efficient], it lacks the context awareness present in transformer-based Natural Language Processing (NLP) models [VaswaniSPUJGKP17]. [mccully1] expanded on this research by examining the application of the BERT and RoBERTa models (both bidirectional) to generate embeddings for lifted code. These embeddings were then used to train neural networks to identify stack-based buffer overflows (CWE-121) within the lifted code. The research introduced in this paper builds on the foundational work of [RN9, schaad2022deeplearningbased, mccully1] by comparing and contrasting the impacts of using a unidirectional transformer (GPT-2), bidirectional transformers (BERT, RoBERTa), and non-transformer-based embedding models (Skip-Gram, Continuous Bag of Words) to train an LSTM vulnerability classifier.
@@ II Methodology and Implementation
Figure 1: Steps to Build GPT-2 Model Using Lifted Code
@@
Figure 2: Steps for Training a Neural Network to Detect Buffer Overflows
@@ II-A Embedding Dataset Creation
The dataset used to train the GPT-2 model was generated by selecting code samples from the NIST SARD Juliet source code dataset without CWE-121 and compiling them on a Kali Linux system [^1] Originally, the samples were also compiled on a Windows 10 system. However, after inspecting the code lifted to LLVM, it was found that RetDec was stripping calls to the standard C library. Thus, these samples were removed from the dataset.. The Juliet dataset contains code samples for several vulnerability classes and the corresponding code samples that have implemented a fix to remediate the vulnerability. The code samples were converted into object files on Linux by running the make Juliet1.3 command.
@@ II-B Neural Network Dataset Creation
A subset of the Juliet dataset was used to create the dataset used to train a neural network to differentiate between LLVM functions that contain buffer overflow vulnerabilities and those that do not. As noted previously, the Juliet dataset contains code samples for several vulnerabilities, including CWE-121 (Stack-Based Buffer Overflow). For this research, these code samples were compiled into object files, and the functions were extracted from these files. The Juliet dataset uses a standard naming scheme that indicates functions that contain buffer overflows and those in which the vulnerability was mitigated. The preprocessing steps in II-C were applied to the data set, but after the preprocessing was completed, it was discovered that some of the functions contained minimal functionality, such as calling a secondary function that contained or did not contain the vulnerability. After preprocessing, most of these functions contained identical instructions and were removed from the dataset when duplicates were eliminated. Finally, functions greater than 2048 tokens in length were removed from the dataset due to resource constraints on neural network training systems. The breakdown of the size of the CWE-121 dataset is given in Table I. Given the relatively small dataset, rather than balancing it with an equal number of vulnerable and non-vulnerable functions, metrics such as precision, recall, and F1-score were utilized to evaluate the overall model performance.
@@ II-C LLVM Preprocessing
The preprocessing phase removes content unique to each function without affecting how the function works. For example, function names are unique to each function but do not affect how the LLVM code executes. Thus, a function name of UniqueFunctionName can be standardized to Func_one without losing functionality. The purpose of preprocessing is to remove these aspects of the LLVM functions to standardize how the functions look. The rationale for this choice is the belief that excessive superfluous information within the LLVM functions could hinder the model’s ability to learn the most relevant code-level patterns. This study followed the preprocessing steps described in detail in [mccully1].
@@ II-D GPT-2 Model Creation
For this study, the GPT2LMHeadModel HuggingFace library was used to create a GPT-2 model that was trained using the same LLVM functions used to train the word2vec, BERT, and RoBERTa models in  [mccully1]. The system used to train the model had dual TI 4090 GPUs. The tokenizer used to tokenize the LLVM functions was the ByteLevelBPETokenizer tokenizer, and the maximum size of the LLVM functions was set to 2048 tokens. The number of hidden layers was set to 12, the dimensionality of the embeddings was set to 100, and the number of attention heads was set to 10. Finally, the embedding dataset detailed in Section II-A was split into two subsets; one subdataset for training that contained 90% of the overall dataset, and a second for validation. The model ran for 20 epochs and steadily improved accuracy every 1,000 steps, as demonstrated in Table II.
@@ IV Discussion
IV-A Unidirectional vs. Bidirectional Embeddings
@@
IV-B Unidirectional vs. Skip-Gram Embeddings
In terms of accuracy, the top-performing neural network using GPT-2 embeddings slightly outperformed the best model with Skip-Gram embeddings. However, the Skip-Gram-based model achieved a significantly higher recall score, while the F1-scores for both models were nearly identical. This suggests that the best models using GPT-2 and Skip-Gram embeddings performed similarly. GPT-2 embeddings consistently achieved accuracies around 92% with the SGD optimizer, a learning rate of 0.0001, and various momentum values. However, models using Skip-Gram embeddings reached this level of accuracy only once, when trained with the Adam optimizer and a learning rate of 0.001 [mccully1].
@@ V-A Dataset Sizes
V-B Data Snooping
@@
*arp2022and [arp2022and] states that data snooping is one of the biggest pitfalls in research on the application of machine learning to problems in the cybersecurity field. One form of data snooping, test snooping, occurs when training and validation data are not strictly segmented. In the context of the current study, test snooping was not a factor since the samples containing CWE-121 were not used to train the GPT-2 model. However, due to the prominence of the issue within existing research [arp2022and], the effects of test snooping on model accuracy were studied in [datasnooping] with no impact found.
V-C Embedding Layer Tuning Enhancements
@@ VI Related Work
Although several papers have employed LLMs to identify vulnerabilities in source code, considerably fewer have focused on using them to detect specific categories of vulnerabilities (e.g., buffer overflows) in compiled code. In [RN67], researchers train a RoBERTa model from scratch using assembly code instructions. The embeddings from this model are used to train a Message Passing Neural Network (MPNN) to differentiate between vulnerable and non-vulnerable code. [gallagherllvm] also built a RoBERTa model to learn code-level semantics to train a neural network to detect vulnerabilities. However, rather than training the RoBERTa model with assembly code, the researchers trained it with LLVM code. [han2022binary] trained a BERT model using P-Code; these P-Code embeddings were used to train a model that could detect vulnerabilities within the lifted code.